Prove that if a sequence {an}∞n=1 satisfies Cauchy’s criterion, then it is bounded


1
Expert's answer
2021-02-03T03:53:01-0500

Let "(a_n)_ {n = 1}^{n = \infty}" be a Cauchy sequence. This implies that given "\epsilon > 0" "\exist" "N \in \mathbb{N}" "\ni" "\mid a_n - a_m \mid < \epsilon" whenever "n,m \geq N" . When "m = N" , we have that "\mid a_n - a_N \mid < \epsilon"

But "\mid a_n \mid - \mid a_N \mid \leq \mid \mid a_n \mid - \mid a_N \mid \mid \leq \mid a_n - a_N \mid < \epsilon"


"\mid a_n\mid < \epsilon + \mid a_N \mid"

Let "K := \max (\mid a_0 \mid, \mid a_1 \mid, \mid a_2 \mid, \cdot \cdot \cdot , \mid a_{N-1} \mid, \mid a_N \mid, \epsilon + \mid a_N \mid )"

So we have that "\mid a_n \mid \leq K" as desired.